OpenStudy (anonymous):

If you have y^2+x^2=40,000 , can you square root the entire side of each equation, so you'd end up with y+x= sqrt (40,000)? If not, how would I solve this equation?

OpenStudy (anonymous):

wrong if you do y+x= sqrt (40,000

OpenStudy (anonymous):

there is infinity solutions

OpenStudy (anonymous):



OpenStudy (anonymous):

Thanks everyone for the answers! However, I'm still a little confused. In the question, I have to find out where two equations cross; in this case, the equation i just mentioned is one of them and I wanted to get it in slope intercept form. So... are you saying that I could simplify the equation in the way I mentioned; I just couldn't solve for it because there is an infinite number of answers?

OpenStudy (anonymous):

can you write both the equations ?

OpenStudy (anonymous):

y=2x+4 and y^2+x^2=40,000

OpenStudy (anonymous):

simply plug y=2x+4 and solve for x (2x+4)^2+x^2=40000 find x and then y from first equation.

OpenStudy (anonymous):

Got it. Thanks! Also, one last quick little clarification for the future so I don't confuse myself: if I had x^2+y^2=40,000, and I just wanted to simplify it, I could square root both sides? I wasn't sure if you were allowed to because it's addition.

OpenStudy (anonymous):

\[4x^2+16x+16+x^2=40,000\] \[5x^2+16x-39984=0\] \[x=\frac{ -16\pm \sqrt{256-4*5*-39984} }{ 2*5 }=\frac{ -16\pm \sqrt{256+799680} }{10 }\] \[=\frac{ -16\pm \sqrt{799936} }{ 10 }=?\]

OpenStudy (anonymous):

\[\sqrt{x^2+y^2+2xy}=x+y\] or \[\left( x+y \right)^2=x^2+y^2+2xy\]

OpenStudy (anonymous):

Okay, got it. Thank you so much for helping!